OpenStudy (henrietepurina):

Select the equations that are parallel and perpendicular to y = -3x - 1 and that pass through the point (3, 1).

OpenStudy (henrietepurina):

parallel: y = -3x + 6 perpendicular: y = 1/3 x + 2 and 2/3 parallel: y = -3x + 10 perpendicular: y = 1/3 x parallel: y = -1/3 x perpendicular: y = -3x parallel: y = 1/3 x + 1 perpendicular: y = 3x - 1

OpenStudy (luigi0210):

Do you know what the slope is for a perpendicular line?

OpenStudy (henrietepurina):

yes

OpenStudy (luigi0210):

Then what is the perpendicular slope for -3?

OpenStudy (henrietepurina):

rise 3 run 1

OpenStudy (henrietepurina):

on oops down 3 right 1

OpenStudy (henrietepurina):



OpenStudy (henrietepurina):

no wait oops again...

OpenStudy (anonymous):

It's not the last two because the slope of a parallel line is the same. The slope of a perpendicular line is the opposite reciprocal of the original slope, so it would be 1/3. The second one passes through (3,1), so that would be it.

OpenStudy (henrietepurina):

Can I trust you? This is my final exam?

OpenStudy (henrietepurina):

what do you think @Luigi0210

OpenStudy (henrietepurina):

can you tell be if b is right?

OpenStudy (henrietepurina):

@Hero @YanaSidlinskiy